# GUARDIAN: Safeguarding LLM Multi-Agent Collaborations with Temporal Graph Modeling

### v1

@@ Abstract
The emergence of large language models (LLMs) enables the development of intelligent agents capable of engaging in complex and multi-turn dialogues. However, multi-agent collaboration face critical safety challenges, such as hallucination amplification and error injection and propagation. This paper presents GUARDIAN, a unified method for detecting and mitigating multiple safety concerns in GUARDing Intelligent Agent collaboratioNs. By modeling the multi-agent collaboration process as a discrete-time temporal attributed graph, GUARDIAN explicitly captures the propagation dynamics of hallucinations and errors. The unsupervised encoder-decoder architecture incorporating an incremental training paradigm, learns to reconstruct node attributes and graph structures from latent embeddings, enabling the identification of anomalous nodes and edges with unparalleled precision. Moreover, we introduce a graph abstraction mechanism based on the Information Bottleneck Theory, which compresses temporal interaction graphs while preserving essential patterns. Extensive experiments demonstrate GUARDIAN’s effectiveness in safeguarding LLM multi-agent collaborations against diverse safety vulnerabilities, achieving state-of-the-art accuracy with efficient resource utilization.
@@ 1 Introduction
The advent of large language models (LLMs) has revolutionized the field of artificial intelligence, enabling the development of intelligent agents capable of engaging in complex, multi-turn dialogues [1, 2, 3]. These LLM-based agents have demonstrated remarkable proficiency in various tasks, from question answering and information retrieval to problem-solving [4, 5, 6]. The multi-agent collaboration systems [7, 8] make multiple LLM agents work together to tackle complex problems through interactive communication and cooperative reasoning [9].
The rise of multi-agent collaboration frameworks, particularly through emerging Agent2Agent (A2A) [10] protocols, presents novel paradigms for AI coordination. While these systems hold immense potential for enhancing AI capabilities, they also introduce unique challenges and safety concerns. These critical safety issues include hallucinations propagating between agents [7, 11, 12, 13, 14, 15], where non-factual information generated by one agent propagates and amplifies through a network of interacting agents. Besides, they also include errors being maliciously introduced and propagated between agents [15, 16, 17, 18, 19, 20], where malicious actors introduce errors, causing previously reliable agents to incorporate and perpetuate these errors. Therefore, we classify these safety issues into two categories: hallucination amplification not involving human factors and error injection and propagation caused by human factors. Injected errors can be further categorized into agent-targeted attacks that compromise agents through malicious system prompts [21] and communication-targeted attacks that corrupt information flow between conversation rounds [19, 22]. Figure 1 illustrates these critical safety issues, demonstrating how hallucinations and injected errors can propagate through a multi-agent system. These problems pose significant risks to LLM-based multi-agent systems, necessitating robust detection and mitigation methods.
Defensive measures against hallucination and error propagation in LLM-based agent collaboration fall into two categories. Collaborative error detection methods, such as cross-examination [23] and external supportive feedback [13], target individual model outputs but overlook propagation dynamics in multi-agent settings. Moreover, multi-agent collaboration approaches employ majority voting strategies [24] or uncertainty estimations [12]. However, they simplify agent dependencies and do not fully capture the complex dependencies in multi-agent systems. Moreover, they require base model modifications, limiting their applicability to close-source models.
@@
This research aims to develop a unified method for detecting and mitigating multiple safety concerns in LLM collaboration processes, including hallucination amplification and error injection and propagation. To achieve this, we model the multi-agent collaboration process into a discrete-time temporal attributed graph, where nodes represent agents at different timesteps, edges represent inter-agent communications and node attributes encode agent responses. This representation allows for explicitly modeling the propagation dynamics of hallucinations and errors through node text attributes while making the collaboration process explicit and transparent. Based on this temporal graph framework, we propose an unsupervised learning paradigm founded on an encoder-decoder architecture. This model learns to reconstruct both node attributes and graph structures from latent embeddings, enabling us to identify anomalies that deviate from normal patterns, precisely pinpointing potential hallucinations or errors with unparalleled precision. To further boost the robustness and efficiency of our framework, we pioneer a graph abstraction mechanism rooted in the Information Bottleneck Theory [25]. This mechanism empowers us to compress temporal interaction graphs by filtering out redundant or irrelevant information while preserving the essential patterns most crucial for anomaly detection. Moreover, our theoretical analysis reveals that information flow between agents remains bounded, with transitions between historical and current states guided by prior constraints.
@@ 2 Related Work
2.1 Safety Problem in Multi-agent Collaborations
The emergence of LLM agents has sparked interest in multi-agent collaboration systems [7, 8, 26], where multiple agents combine capabilities to tackle complex tasks [1, 9]. Recent research has highlighted critical safety challenges in LLM multi-agent collaborations, particularly regarding hallucination amplification and error injection and propagation. DebUnc [12] pioneered the study of uncertainty in multi-agent debates, demonstrating how hallucinations spread through interactions. MultiAgent Collaboration Attack [18] and Huang et al. [27] reveal how adversarial agents exploit collaborative dynamics to manipulate outcomes, highlighting the need for robust defense mechanisms.
2.2 Defenses in Multi-agent Collaborations
Current defenses against hallucination and error injection and propagation in LLM-based agent collaboration fall into two categories. The first approach focuses on detecting factual errors through cross-examination [23] and external feedback [13], but fails to model propagation dynamics in multi-agent settings. The second approach targets multi-agent collaboration directly through majority voting mechanisms [24], yet relies on oversimplified assumptions about agent independence.
@@ 3.1 Safety Problems in LLM Agents Collaboration
  • Communication-targeted attacks: During time $t-1$ to $t$, an adversary corrupts the communication from $m_{i}$ to $m_{j}$, where $m_{j}$ receives a modified $r_{t-1,i}$ at time $t$.
@@ 3.2 Temporal Attributed Graph Representation
Message passing. Messages flow through network nodes $\mathcal{V}$ and edges $\mathcal{E}$, reflecting LLM discussions. Forward propagation across timesteps produces output $o$ given query $q$.
@@
Figure 3: Framework overview of GUARDIAN, showing a case study at timestep $t_{2}$. (1) Graph Preprocessing: The collaboration information from $t_{0}$ to $t_{2}$ is transformed into node attributes $\bm{x}_{t,i}$ and graph structures $\mathcal{E}$ using BERT and communication pattern abstraction. (2) Attributed Graph Encoder processes each time’s graph to obtain node embeddings $\{\mathbf{Z}_{t}\}_{t=1}^{T}$. (3) Time Information Encoder aggregates multi-timestamp graph embeddings into the final timestamp $\mathbf{Z}_{T}$. (4) Structure and Attribute Reconstruction Decoder output reconstructed graph ^$\hat{\mathcal{E}}_{T}$ and node attributes $\hat{\mathbf{X}}_{T}$. (5) Anomaly scores $s_{v}$, calculated from the original and reconstructed graphs, identify and exclude the highest-scoring anomalous node from subsequent iterations.
@@ 4.1 Problem Formulation
[…]
@@
where $s_{v}\in f(\mathcal{G}_{t})$, and we further obtain the updated graph sequence $\mathcal{G}_{t}^{\prime}=(\mathcal{V}_{t}\setminus\mathcal{V}_{t}^{*},\mathcal{E}_{t}\setminus\mathcal{E}_{\mathcal{V}_{t}^{*}})$. To tackle this optimization problem, we focus on the carefully designed loss function $\mathcal{L}(f)$ that guides the anomaly detection model, and robust scoring mechanisms that compute node-level anomaly scores $s_{v}$. These scores enable the identification of anomalous nodes $\mathcal{V}_{t}^{*}$ through a reconstruction-based approach. The cleaned graph $\mathcal{G}_{t}^{\prime}$ is obtained by removing the identified anomalous elements.
@@ 4.2 Encoder-Decoder Architecture
Attributed graph encoder. In multi-agent collaboration, node attributes and inter-agent connections are essential information carriers. At time step $t$, we employ GCN to learn node embeddings $\mathbf{Z}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|\times d}$ ($d$ is embedding dimension) by aggregating information from neighboring nodes. Given node feature $\mathbf{X}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|\times k}$ ($k$ is feature dimension) and structure $\mathcal{E}_{t}$, GCN compresses node attributes and neighborhood information through the layer-wise propagation rule $\mathbf{H}^{(l+1)}=ReLU(\mathbf{D}_{e}^{-\frac{1}{2}}\mathbf{A}_{e}\mathbf{D}_{e}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)})$, where $\mathbf{H}^{(0)}=\mathbf{X}_{t}$. By stacking two GCN layers, nodes can perceive 2-hop neighborhood information while maintaining computational efficiency.
Time information encoder. In multi-agent systems, agents’ current decisions depend on historical interaction patterns. To capture these temporal dependencies, we employ a Transformer-based temporal encoder to process the sequence of graph embeddings $\{\mathbf{Z}_{1},\mathbf{Z}_{2},...,\mathbf{Z}_{T}\}$ obtained from the GCN encoder. Through the core self-attention mechanism $Attn(\mathbf{Q},\mathbf{K},\mathbf{V})=softmax(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}})\mathbf{V}$, where $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ are query, key, and value matrices derived from the input sequence, the model dynamically weighs and aggregates information from previous interaction rounds, focusing on the most relevant historical patterns when generating the final representation $\mathbf{Z}_{T}$.
@@ 4.4 Incremental Training Paradigm
Table 1: Accuracy (%) comparison of GPT-3.5-turbo, GPT-4o and Claude-3.5-sonnet under hallucination amplification or two types of error injection and propagation.
@@ 5.1 Experimental Settings
Datasets. Our evaluation employs four benchmark datasets that span diverse domains and cognitive requirements. The benchmark datasets include MMLU [36], MATH [37], FEVER [38], and Biographies [1]. Following [9, 12], we randomly sample 100 questions from each dataset, with three independent testing iterations to ensure statistical robustness.
Compared methods. We evaluate against three baseline categories: (1) standard multi-agent frameworks without defense (LLM Debate [1], DyLAN [9]); (2) hallucination detection methods (SelfCheckGPT [39]); and (3) error detection approaches (Challenger [27], Inspector [27]). LLM Debate serves as the fundamental framework, with several others building upon it. Details are in Appendix A.3. We present two variants: GUARDIAN.s performs anomaly detection using static graphs from the current time-step, while GUARDIAN incorporates historical graph information.
Implementation details. We evaluate models in a zero-shot CoT [40] setting using both closed-sourced and open-sourced models (GPT-3.5-turbo [41], GPT-4o [42], Claude-3.5-sonnet [43], Llama3.1-8B [44]), with all agents treated without role differentiation. We primarily test with 4 agents, conducting additional experiments with 3-7 agents, with detailed prompts in Appendix A.4.
Evaluation metrics. We evaluate both model accuracy, anomaly detection rate and API calls. For error injection and propagation, we simulate errors through agent-targeted attacks (randomly selecting 1 agent in the first round) and communication-targeted attacks (randomly interfering with multiple communication edges in intermediate rounds), simulating real-world malicious agents and communication interference, respectively. The detailed experimental setup is provided in Appendix A.5.
@@ 5.2 Experimental Results
Our model’s anomaly detection rate results, detailed in Appendix A.6, average above 80% with a peak of 94.74%. Figure 4 shows a real case from the MATH dataset containing both hallucination and agent-targeted errors. In the first round, the four nodes fail to reach consensus, and GUARDIAN removes Node 1 as an anomaly due to its highest score. The remaining three nodes still fail to reach consensus in the second round, leading to the removal of Node 3. The final round between the two remaining nodes achieves consensus with the correct answer of 8. This process validates the effectiveness, as Node 1 was indeed a malicious agent and Node 3 exhibited hallucination.
Scalability analysis. To evaluate scalability, we varied the number of agents from 3 to 7 using GPT-3.5-turbo. As shown in Table 6, our framework maintains consistent performance across different configurations, demonstrating effective adaptation to multi-agent networks of varying sizes. This robust scalability stems from our temporal graph modeling approach, which efficiently captures agent interactions regardless of network size.
Running cost. As shown in Figure 5, our approach achieves optimal performance with the lowest API calls among all baselines. Our incremental node pruning strategy both removes anomalous nodes and naturally reduces redundant API queries, which is effective as agents tend to reach consensus in multi-round debates. Additional experimental results are provided in Appendix A.7,  A.8, and A.9.

### v2

@@ Abstract
The emergence of large language models (LLMs) enables the development of intelligent agents capable of engaging in complex and multi-turn dialogues. However, multi-agent collaboration faces critical safety challenges, such as hallucination amplification and error injection and propagation. This paper presents GUARDIAN, a unified method for detecting and mitigating multiple safety concerns in GUARDing Intelligent Agent collaboratioNs. By modeling the multi-agent collaboration process as a discrete-time temporal attributed graph, GUARDIAN explicitly captures the propagation dynamics of hallucinations and errors. The unsupervised encoder-decoder architecture incorporating an incremental training paradigm learns to reconstruct node attributes and graph structures from latent embeddings, enabling the identification of anomalous nodes and edges with unparalleled precision. Moreover, we introduce a graph abstraction mechanism based on the Information Bottleneck Theory, which compresses temporal interaction graphs while preserving essential patterns. Extensive experiments demonstrate GUARDIAN’s effectiveness in safeguarding LLM multi-agent collaborations against diverse safety vulnerabilities, achieving state-of-the-art accuracy with efficient resource utilization. The code is available at https://github.com/JialongZhou666/GUARDIAN.
@@ 1 Introduction
The advent of large language models (LLMs) has revolutionized the field of artificial intelligence, enabling the development of intelligent agents capable of engaging in complex, multi-turn dialogues [1, 2, 3, 4, 5]. These LLM-based agents have demonstrated remarkable proficiency in various tasks, from question answering and information retrieval to problem-solving [6, 7, 8]. The multi-agent collaboration systems [9, 10] make multiple LLM agents work together to tackle complex problems through interactive communication and cooperative reasoning [11].
The rise of multi-agent collaboration frameworks, particularly through emerging Agent2Agent (A2A) [12] protocols, presents novel paradigms for AI coordination. While these systems hold immense potential for enhancing AI capabilities, they also introduce unique challenges and safety concerns. These critical safety issues include hallucinations propagating between agents [9, 13, 14, 15, 16, 17, 18], where non-factual information generated by one agent propagates and amplifies through a network of interacting agents. Besides, they also include errors being maliciously introduced and propagated between agents [17, 19, 20, 21, 22, 23], where malicious actors introduce errors, causing previously reliable agents to incorporate and perpetuate these errors. Therefore, we classify these safety issues into two categories: hallucination amplification not involving human factors and error injection and propagation caused by human factors. Injected errors can be further categorized into agent-targeted attacks that compromise agents through malicious system prompts [24] and communication-targeted attacks that corrupt information flow between conversation rounds [22, 25]. Figure 1 illustrates these critical safety issues, demonstrating how hallucinations and injected errors can propagate through a multi-agent system. These problems pose significant risks to LLM-based multi-agent systems, necessitating robust detection and mitigation methods.
Defensive measures against hallucination and error propagation in LLM-based agent collaboration fall into two categories. Collaborative error detection methods, such as cross-examination [26] and external supportive feedback [15], target individual model outputs but overlook propagation dynamics in multi-agent settings. Moreover, multi-agent collaboration approaches employ majority voting strategies [27] or uncertainty estimations [14]. However, they simplify agent dependencies and do not fully capture the complex dependencies in multi-agent systems. Moreover, they require base model modifications, limiting their applicability to closed-source models.
@@
This research aims to develop a unified method for detecting and mitigating multiple safety concerns in LLM collaboration processes, including hallucination amplification and error injection and propagation. To achieve this, we model the multi-agent collaboration process into a discrete-time temporal attributed graph, where nodes represent agents at different timesteps, edges represent inter-agent communications and node attributes encode agent responses. This representation allows for explicitly modeling the propagation dynamics of hallucinations and errors through node text attributes while making the collaboration process explicit and transparent. Based on this temporal graph framework, we propose an unsupervised learning paradigm founded on an encoder-decoder architecture. This model learns to reconstruct both node attributes and graph structures from latent embeddings, enabling us to identify anomalies that deviate from normal patterns, precisely pinpointing potential hallucinations or errors with unparalleled precision. To further enhance the robustness and efficiency of our framework, we develop a graph abstraction mechanism grounded in Information Bottleneck Theory [28], which represents, to the best of our knowledge, the first systematic application of IB principles to mitigate safety risks in LLM-based multi-agent collaboration. This mechanism empowers us to compress temporal interaction graphs by filtering out redundant or irrelevant information while preserving the essential patterns most crucial for anomaly detection. Moreover, our theoretical analysis reveals that information flow between agents remains bounded, with transitions between historical and current states guided by prior constraints.
@@ 2 Related Work
Safety problem in multi-agent collaborations. The emergence of LLM agents has sparked interest in multi-agent collaboration systems [9, 10, 29], where multiple agents combine capabilities to tackle complex tasks [1, 11]. Recent research has highlighted critical safety challenges in LLM multi-agent collaborations, particularly regarding hallucination amplification and error injection and propagation. DebUnc [14] pioneered the study of uncertainty in multi-agent debates, demonstrating how hallucinations spread through interactions. MultiAgent Collaboration Attack [21] and Huang et al. [30] reveal how adversarial agents exploit collaborative dynamics to manipulate outcomes, highlighting the need for robust defense mechanisms.
Defenses in multi-agent collaborations. Current defenses against hallucination and error injection and propagation in LLM-based agent collaboration fall into two categories. The first approach focuses on detecting factual errors through cross-examination [26] and external feedback [15], but fails to model propagation dynamics in multi-agent settings. The second approach targets multi-agent collaboration directly through majority voting mechanisms [27], yet relies on oversimplified assumptions about agent independence.
@@ 3.1 Safety Problems in LLM Agents Collaboration
  • Communication-targeted attacks: During time t1t-1 to $t$, an adversary corrupts the communication from $m_{i}$ to $m_{j}$, where $m_{j}$ receives a modified $r_{t-1,i}$ at time $t$.
@@ 3.2 Temporal Attributed Graph Representation
Message passing. Messages flow through network nodes $\mathcal{V}$ and edges $\mathcal{E}$, reflecting LLM discussions. Forward propagation across timesteps produces output $o$ given query $q$. Agents in each round can only access outputs from directly connected agents in the immediately preceding round, reflecting realistic communication constraints in distributed systems.
@@
Figure 3: Framework overview of GUARDIAN, showing a case study at timestep $t_{2}$. (1) Graph Preprocessing: The collaboration information from $t_{0}$ to $t_{2}$ is transformed into node attributes $\bm{x}_{t,i}$ and graph structures $\mathcal{E}$ using BERT and communication pattern abstraction. (2) Attributed Graph Encoder processes each time’s graph to obtain node embeddings $\{\mathbf{Z}_{t}\}_{t=1}^{T}$. (3) Time Information Encoder aggregates multi-timestamp graph embeddings into the final timestamp $\mathbf{Z}_{T}$. (4) Structure and Attribute Reconstruction Decoder output reconstructed graph $\hat{\mathcal{E}}_{T}$ and node attributes $\hat{\mathbf{X}}_{T}$. (5) Anomaly scores $s_{v}$, calculated from the original and reconstructed graphs, identify and exclude the highest-scoring anomalous node from subsequent iterations.
@@ 4.1 Problem Formulation
$\min_{f}\mathcal{L}(f)\quad\text{s.t.}$ $f:\mathcal{G}_{t}\rightarrow\mathbb{R}^{|\mathcal{V}_{t}|+|\mathcal{E}_{t}|},$ (1)
@@
where $s_{v}\in f(\mathcal{G}_{t})$, and $\mathcal{L}(f)$ denotes the loss function used to train the anomaly detection model. We further obtain the updated graph sequence $\mathcal{G}_{t}^{\prime}=(\mathcal{V}_{t}\setminus\mathcal{V}_{t}^{*},\mathcal{E}_{t}\setminus\mathcal{E}_{\mathcal{V}_{t}^{*}})$. To tackle this optimization problem, we focus on the carefully designed loss function $\mathcal{L}(f)$ that guides the anomaly detection model, and robust scoring mechanisms that compute node-level anomaly scores $s_{v}$. These scores enable the identification of anomalous nodes $\mathcal{V}_{t}^{*}$ through a reconstruction-based approach. The cleaned graph $\mathcal{G}_{t}^{\prime}$ is obtained by removing the identified anomalous elements.
@@ 4.2 Encoder-Decoder Architecture
Attributed graph encoder. In multi-agent collaboration, node attributes and inter-agent connections are essential information carriers. At time step $t$, we employ GCN to learn node embeddings $\mathbf{Z}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|\times d}$ ($d$ is the embedding dimension) by aggregating information from neighboring nodes. Given node feature $\mathbf{X}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|\times k}$ ($k$ is feature dimension) and structure $\mathcal{E}_{t}$, GCN compresses node attributes and neighborhood information through the layer-wise propagation rule $\mathbf{H}^{(l+1)}=ReLU(\mathbf{D}_{e}^{-\frac{1}{2}}\mathbf{A}_{e}\mathbf{D}_{e}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)})$, where $\mathbf{H}^{(0)}=\mathbf{X}_{t}$. By stacking two GCN layers, nodes can perceive 2-hop neighborhood information while maintaining computational efficiency.
Time information encoder. In multi-agent systems, agents’ current decisions depend on historical interaction patterns. To capture these temporal dependencies, we employ a Transformer-based temporal encoder to process the sequence of graph embeddings $\{\mathbf{Z}_{1},\mathbf{Z}_{2},...,\mathbf{Z}_{T}\}$ obtained from the GCN encoder. Through the core self-attention mechanism $Attn(\mathbf{Q},\mathbf{K},\mathbf{V})=softmax(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}})\mathbf{V}$, where $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ are query, key, and value matrices derived from the input sequence, the model dynamically weighs and aggregates information from previous interaction rounds, focusing on the most relevant historical patterns when generating the final representation $\mathbf{Z}_{T}$. This approach implicitly models temporal dependencies between nodes across different timesteps through cross-time self-attention, enabling the capture of how agent interactions evolve and influence each other over time.
@@ 4.4 Incremental Training Paradigm
Table 1: Accuracy (%) comparison of GPT-3.5-turbo, GPT-4o and Claude-3.5-sonnet under hallucination amplification or two types of error injection and propagation. Bold values represent the highest accuracy.
@@ 5.1 Experimental Settings
Datasets. Our evaluation employs four benchmark datasets that span diverse domains and cognitive requirements. The benchmark datasets include MMLU [39], MATH [40], FEVER [41], and Biographies [1]. Specifically, these benchmarks test broad multi-subject knowledge (MMLU), mathematical reasoning (MATH), and factual verification against evidence (FEVER, Biographies). Following [11, 14], we randomly sample 100 questions from each dataset, with three independent testing iterations to ensure statistical robustness. We train a separate model for each dataset and apply the Incremental Training Paradigm within each dataset, focusing on in-distribution anomaly detection where the data distribution remains consistent across episodes.
Compared methods. We evaluate against three baseline categories: (1) standard multi-agent frameworks without defense (LLM Debate [1], DyLAN [11]); (2) hallucination detection methods (SelfCheckGPT [42]); and (3) error detection approaches (Challenger [30], Inspector [30]). This allows us to compare against both the inherent resilience of standard debate-based frameworks and dedicated defense strategies that employ self-consistency checks or deploy specialized inspector agents. LLM Debate serves as the fundamental framework, with several others building upon it. Details are in Appendix A.3. We present two variants: GUARDIAN.s performs anomaly detection using static graphs from the current time step, while GUARDIAN incorporates historical graph information.
Implementation details. We evaluate models in a zero-shot CoT [43] setting using both closed-sourced and open-sourced models (GPT-3.5-turbo [44], GPT-4o [45], Claude-3.5-sonnet [46], Llama3.1-8B [47]), with all agents treated without role differentiation. We primarily test with 4 agents, conducting additional experiments with 3-7 agents. Runtime efficiency is evaluated under communication-targeted attacks using consistent experimental settings. Detailed prompts are provided in Appendix A.4.
Evaluation metrics. We evaluate model accuracy, anomaly detection rate, False Discovery Rate (FDR), API calls, and runtime efficiency. For error injection and propagation, we simulate errors through agent-targeted attacks (randomly selecting 1 agent in the first round) and communication-targeted attacks (randomly interfering with multiple communication edges in intermediate rounds), simulating real-world malicious agents and communication interference, respectively. While the above tests utilize fully connected multi-agent topologies, we also conduct experiments with sparse communication graphs to validate our method’s effectiveness across different network structures, where each agent connects to only 25%, 50%, or 75% of agents in subsequent rounds. The detailed experimental setup is provided in Appendix A.5.
@@ 5.2 Experimental Results
Table 2: Analysis of Detection Reliability: False Discovery Rate (FDR, %).
Safety Issue Datasets
MMLU MATH FEVER Biographies
Hallucination Amplification 26.67 13.11 17.86 15.69
Agent-targeted Error Injection and Propagation 22.22 8.32 20.53 13.23
Communication-targeted Error Injection and Propagation 30.67 18.42 28.57 19.65
@@
Our model’s anomaly detection rate results, detailed in Appendix A.6, average above 80% with a peak of 94.74%. We also analyzed potential failure modes where GUARDIAN may falsely remove correct agents. Table 2 presents False Discovery Rate (FDR) results with GPT-3.5-turbo across 4 agents, demonstrating consistently low rates across different scenarios, with an exceptionally low rate of 8.32% on the MATH dataset under agent-targeted attacks and remaining well below 20% for most scenarios. Additional analysis reveals that even when correct agents are mistakenly removed, the impact on overall system performance is minimal due to natural redundancy in multi-agent collaborations and GUARDIAN’s conservative strategy of removing only one agent per round, ensuring graceful degradation rather than catastrophic error amplification. Figure 4 shows a real case from the MATH dataset containing both hallucination and agent-targeted errors. In the first round, the four nodes fail to reach consensus, and GUARDIAN removes Node 1 as an anomaly due to its highest score. The remaining three nodes still fail to reach consensus in the second round, leading to the removal of Node 3. The final round between the two remaining nodes achieves consensus with the correct answer of 8. This process validates the effectiveness, as Node 1 was indeed a malicious agent and Node 3 exhibited hallucination.
Topology robustness. To further validate our approach’s robustness across different network topologies, Table 5.2 shows performance under sparse communication graphs in hallucination amplification scenarios, where agents connect to only 25%, 50%, or 75% of other agents in subsequent rounds. GUARDIAN consistently outperforms baselines across all sparsity levels, demonstrating effective performance beyond fully connected scenarios and validating our structure reconstruction decoder’s utility in non-trivial topologies.
Scalability analysis. To evaluate scalability, we varied the number of agents from 3 to 7 using GPT-3.5-turbo. As shown in Table 5.2, our framework maintains consistent performance across different configurations, demonstrating effective adaptation to multi-agent networks of varying sizes. This robust scalability stems from our temporal graph modeling approach, which efficiently captures agent interactions regardless of network size.
Running cost. As shown in Figure 5, our approach achieves optimal performance with the lowest API calls among all baselines. Our incremental node pruning strategy both removes anomalous nodes and naturally reduces redundant API queries, which is effective as agents tend to reach consensus in multi-round debates. Regarding runtime efficiency, Table 6 demonstrates that our method achieves the lowest average runtime per question on MMLU and FEVER datasets. On MATH, it incurs only marginal overhead (less than 5 seconds on average) compared to the baseline LLM Debate, while remaining significantly faster than other defense-enabled methods. The reduced runtime stems from our node deletion strategy, which decreases both API calls and inter-agent waiting time per round. Although anomaly detection is applied in each round, it introduces minimal computational overhead due to the small scale of agent communication graphs, contributing only a negligible fraction to the total runtime. Additional experimental results are provided in Appendix A.7,  A.8, and A.9.
Table 3: Accuracy (%) comparison under different connection sparsity under hallucination amplification on MATH dataset with GPT-3.5-turbo. Bold values represent the highest accuracy.
Method Sparsity
25% 50% 75%
LLM Debate 26.5 34.6 34.2
DyLAN 38.6 41.6 39.3
SelfCheckGPT 7.6 5.6 3.4
GUARDIAN 52.2 56.1 57.3
@@
Table 5: Runtime cost (s) comparison under communication-targeted attacks with GPT-3.5-turbo. Bold values represent the lowest time cost.
Method MMLU MATH FEVER
LLM Debate 29.26 40.31 25.02
Challenger 26.15 56.23 27.5
Inspector 76.82 129.59 69.25
GUARDIAN 18.89 45.19 17.13